OpenStudy (icalibear):

Did I write this equation in correct slope-intercept form? What is an equation for the line that passes through the points (-1, -4) and (1, 4)? Write the equation in slope-intercept form. My answer: y = 4/1x + 4

OpenStudy (chihiroasleaf):

is your answer y = (4/1 ) x + 4 ? it's already in slope-intercept form, but I think it's not correct can you show how do you solve this problem? :)

OpenStudy (icalibear):

Well it told me to write an equation for the line that passes through the points (-1, -4) and (1, 4) so I did this: slope-intercept form: y = mx + b slope = m = rise/run = m = y2-y1/x2-x1 plug it in: 4 - -4/1 - -1 = 8/2 = 4/1 y-intercept is 4 given the data plug it in: y = 1/4x + 4

OpenStudy (icalibear):

@chihiroasleaf

OpenStudy (icalibear):

4/1x** not 1/4x

OpenStudy (chihiroasleaf):

hhmm.., I see.., you've found the slope correctly :) but, in slope-intercept form, the 'intercept' means the y-intercept of the line, so the point in which the line cross the y-axis, so the coordinate of x is 0, and it's not given in the problem, right? to find the equation of line, if you know the slope and one point on the line, let the point is (x1,y1), you can use the formula y - y1 = m ( x - x1) or if you know two points on the line (x1,y1) and (x2, y2), you can use formula \[\frac{ y - y _{1} }{ y _{2} - y _{1}} = \frac{ x - x_{1} }{ x_{2} - x_{1} }\] then simplify the equation into slope-intercept form... :) can you continue it?

OpenStudy (icalibear):

so it would be y - -4/4 - -4, which is y/8

OpenStudy (icalibear):

and then x - -1/1--1 which is x/8

OpenStudy (icalibear):

ahh i dont know (/.\)

OpenStudy (chihiroasleaf):

 can you continue it? :)

OpenStudy (icalibear):

Well the final fraction would be 4/1, but I don't know how to get the y intercept